# Yet another characterization of Birkhoff Billiards inside discs

### 17 Mar 2026 v1

@@ 1. Introduction and statement of our main results
1.1. Comments and ideas for future developments
@@ 2.1. Angular functions
2.2. Relation between angular functions and high-order Lazutkin coordinates
@@
Since $Z_{\El,0}=\text{Id}$ and we are in Lazutkin parametrization, we conclude that $Z_{\El,0}$ is an angular function of order $1$, hence $\Phi^{+}_{2}=0$ (this follows also from an explicit computation).
@@
Carrying out the algebra yields the desired formula. ∎ We conclude that, in order for $Z$ to be an angular function of order $2$, the function $A=Z_{\El,1}$ solves the following differential equation:
@@
Among the four terms involving $\rho$, three have differential degree $3$ and the remaining term has differential degree $1$ (once renormalized with $C^{2}\rho^{2/3}$). Terms with different differential degrees come from different terms in the expansion of the complex exponential function in (2.7).
@@
where $\tilde{\beta}_{4}$ is defined in Remark 0.6. This is of course to be expected, since $\tilde{\beta}_{4}$ is the correction of order $4$ to the angular component $y$ of the billiard map in classical Lazutkin coordinates, and $A$ is exactly designed to compensate such correction.
@@ 3. Proof of the Main Theorem
where $C_{\textrm{BV}}$ is an integration constant that can be fixed in such a way to ensure periodicity of $A$ — that is, we set $\int A^{\prime}=0$, which yields, after integrating by parts, the manifestly positive quantity:
@@
This perhaps surprising relation is of course not accidental. As remarked in Remark 0.7, the function $A$ appears (in some form) in the asymptotic distribution of collision points of periodic orbits; this in turn is tightly related with the asymptotics of the length of periodic orbits of small rotation number –or, equivalently, the coefficients of the Mather’s $\beta$ function at 0– and it is well known (see once again [12, Theorem 5.15] and [15]) that such coefficients are algebraically related with the Marvizi–Melrose invariants.

### 18 Mar 2026 v2

@@ 1. Introduction and statement of our main results
$1.$ Comments and ideas for future developments
@@ 2.1. Angular functions
$2.$ Relation between angular functions and high-order Lazutkin coordinates
@@
Since $Z_{\El,0}=\text{Id}$ and we are in Lazutkin parametrization, we conclude that $Z_{\El,0}$ is an angular function of order 11, hence $\Phi^{+}_{2}=0$ (this follows also from an explicit computation).
@@
Carrying out the algebra yields the desired formula. ∎ We conclude that, in order for $Z$ to be an angular function of order 22, the function $A=Z_{\El,1}$ solves the following differential equation:
@@
Among the four terms involving $\rho$, three have differential degree 33 and the remaining term has differential degree 11 (once renormalized with $C^{2}\rho^{2/3}$). Terms with different differential degrees come from different terms in the expansion of the complex exponential function in (2.7).
@@
where $\tilde{\beta}_{4}$ is defined in Remark 0.6. This is of course to be expected, since $\tilde{\beta}_{4}$ is the correction of order 44 to the angular component $y$ of the billiard map in classical Lazutkin coordinates, and AA is exactly designed to compensate such correction.
@@ 3. Proof of the Main Theorem
where $C_{\textrm{BV}}$ is an integration constant that can be fixed in such a way to ensure periodicity of AA — that is, we set $\int A^{\prime}=0$, which yields, after integrating by parts, the manifestly positive quantity:
@@
This perhaps surprising relation is of course not accidental. As remarked in Remark 0.7, the function AA appears (in some form) in the asymptotic distribution of collision points of periodic orbits; this in turn is tightly related with the asymptotics of the length of periodic orbits of small rotation number –or, equivalently, the coefficients of the Mather’s $\beta$ function at 0– and it is well known (see once again [12, Theorem 5.15] and [15]) that such coefficients are algebraically related with the Marvizi–Melrose invariants.